Circular No. 7804

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002aj, 2002ak, 2002al, 2002am A. Becker, T. Tyson, and D. Wittman, Bell Laboratories; M. Lopez-

Caniego, Universidad Autonoma de Madrid; D. Norman, Cerro Tololo Interamerican Observatory (CTIO); A. Clocchiatti, Universidad Católica de Chile; I. Dell’Antonio, Brown University; and G. Squires, Space Infrared Telescope Facility, California Institute of Technology, on behalf of the Deep Lens Survey team, report the detection of four apparent supernovae in V - band images taken with the CTIO 4-m reflector (+ MOSAIC imager).

SN  2002 UT  α_{2000}  δ_{2000}  V  Offset
2002aj Jan. 11.269 10^{h}48^{m}56.^{s}56 − 5^{o}32′14′′.1 23.1 1′′.5 W, 1′′.8 S
2002ak Jan. 11.269 10 49 04.32 − 5 55 33.7 22.6  ?
2002al Jan. 11.311 10 50 21.71 − 4 15 05.9 23.6 3′′.1 E
2002am Jan. 16.3  14 02 03.14 −10 35 23.9 19.1 2′′.9 W, 7′′.4 S

SNe 2002aj and 2002ak were not present in images taken on 2001 Mar. 25.12 UT, and SN 2002al was not present on 2001 Mar. 27.04 (all 2001 images involve five stacked 600-s V -band frames, with limiting mag V = 25.5). Rband images of SN 2002ak show a faint, diffuse host galaxy. SN 2002am was found by subtracting images taken on 2001 Apr. 24.3 from images obtained on 2002 Jan. 16.3; it was confirmed with additional images and a spectrum taken by A. Clocchiatti on Jan. 18.3 with the Las Campanas Observatory 6.5-m Baade telescope (+ dual imager/spectrograph LDSS2). The fully reduced spectrum (range 400–800 nm; resolution 1.2 nm) shows that SN 2002am is a type-Ia supernova, some 100 days after maximum light. The approximate redshift of the parent galaxy is 0.03, obtained by matching the features in the supernova spectrum with those of SN 1989B, 96 days after maximum light. Additional photometry: SN 2002aj, 2002 Jan. 14.273, R = 22.3; 14.285, V = 22.9. SN 2002ak, Jan. 14.273, R = 22.7; 14.285, V = 23.3. SN 2002al, Jan. 14.220, R = 22.8; 14.237, V = 23.9.

SUPERNOVA 2001ig IN NGC 7424 Unfiltered CCD magnitudes by C. Bembrick, Bathurst, N.S.W.: 2001

Dec. 19.419 UT, 13.75; 21.435, 13.51; 22.435, 13.40; 24.433, 13.24; 26.430, 13.12; 27.443, 13.05; 28.4347, 13.00; 30.444, 12.96; 2002 Jan. 1.432, 13.00; 3.436, 12.94; 7.4451, 13.02. Visual magnitude estimates: Jan. 6.540, 12.3 (A. Pearce, Nedlands, W. Australia); 11.45, 12.8 (R. Evans, Hazelbrook, N.S.W.); 16.45, 13.0 (Evans); 18.537, 12.9 (Pearce).

2002 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green